(a) x 2 hours (c) x- 12 hours b) x2 5 hours (d)x4 5 hours slop. to tree deinal places as needed Round the yelecept to two decimal places扔needed } Ос. O 0. a.
48 (c)x 12 hours d)x-4 5 hours O A. 623 O B. 51.1 O C. 484 O D. not meaninghul (b) Predict the value of y for x 25 Choose the correct answer below O A. 623 B 1040 ОС. 51.1 D. not meaningh (c) Predict the value of y for x 12 Choose the correct answer bellow O A. 51.1 8
(c) Predict the value of y for x 12. Choose the correct answer belovw OA. 51.1 O B. 484 O C. 104.0 O D. not meaningful OA 623 B. 484 С. 1040 O D. not meaningful O Type here to search 2
Answer #1
X Y XY X^2 Y^2
0 39 0 0 1521
1 44 44 1 1936
3 52 156 9 2704
3 48 144 9 2304
4 63 252 16 3969
6 72 432 36 5184

b X PREDICTED Y = a - b2x

From the above table and formual we get the value are as below:

n 6
sum(XY) 1028.00
sum(X) 17.00
sum(Y) 318.00
sum(X^2) 71.00
sum(Y^2) 17618.00
b 5.5620
a 37.2409

y = a + bx

= 37.24 + 5.562x

a)

when x = 2

y = 37.24 + 5.562x
= 37.24 + 5.562 * 2
= 48.364


b)
when x = 2.5
y = 37.24 + 5.562x
= 37.24 + 5.562 * 2.5
= 51.145

c)
when x = 12

y = 37.24 + 5.562x
= 37.24 + 5.562 * 12
= 103.984


d)

when x = 4.5
y = 37.24 + 5.562x
= 37.24 + 5.562 * 4.5
= 62.269